### v2

# Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases

Abstract
Topological phase transitions beyond anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) and double semion (DS) phases, which has two distinct $\mathbb{Z}_{2}$ topological orders in (2+1)(2+1)D. Previous studies reveal that the transition between them can be either first order or via an intermediate phase, thus the existence of a directly continuous transition between them remains a long-standing problem. Motivated by the fact that both phases can arise from condensing distinct anyons in the $\mathbb{Z}_{4}$ topological order, we introduce a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model to study the TC-DS transition. We confirm the existence of a continuous (2+1)(2+1)D XY* transition between the TC and DS phases by mapping it to a two-coupled quantum Ising model. Importantly, using the condensation order parameters and the area law coefficients of the Wilson loops, we further reveal that $\mathbb{Z}_{4}$ anyons, fractionalized from the $\mathbb{Z}_{2}$ topological orders, become deconfined at the transition between $\mathbb{Z}_{2}$ topologically ordered phases. Our results open a path toward developing a theoretical framework for topological phase transitions beyond anyon condensation.
Introduction. Topological phases of matter, first discovered in the fractional quantum Hall effect [1, 2], have revolutionized our understanding of quantum many-body systems. Owing to their potential applications in quantum error correction and fault-tolerant quantum computing [3, 4], topological phases have become a central focus of both theoretical and experimental research. Exactly solvable models with topologically ordered ground states have been constructed to study these phases [4, 5, 6], and some have been experimentally realized using quantum processors and quantum simulators [7, 8, 9, 10]. As phases of matter beyond the conventional Ginzburg-Landau-Wilson (GLW) paradigm of spontaneous symmetry breaking, topologically ordered states pose significant challenges for studying their phase transitions within the framework of the GLW paradigm. Nonetheless, topological phase transitions driven by anyon condensation can be partially understood using the GLW paradigm, because they can be mapped, via duality transformations, to spontaneous symmetry-breaking phase transitions characterized by local order parameters [11, 12, 13, 14, 15, 16], and such topological phase transitions have been extensively studied using various models and methods [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 14, 15, 27, 28].
Figure 1: Schematic illustration of the motivation and main results. (a) Previous works [29, 30, 31] reveal that the transition between the TC and DS phases can be first order or via an intermediate phase. The existence of a direct continuous transition between them is an open question. (b) We confirm a direct continuous TC-DS transition point, where the adjacent TC and DS phases spontaneously break two different 1-form symmetries. We further reveal deconfined $\mathbb{Z}_{4}$ anyons at their transition point.
However, topological phase transitions that are not induced by anyon condensation remain poorly understood. A notable example is the transition between the toric code (TC) [4] and double semion (DS) phases [6], which represent two distinct $\mathbb{Z}_{2}$ topological orders with incompatible anyons. By interpolating fixed TC and DS string-net models using a one-parameter Hamiltonian, the transition has been studied using exact diagonalization [29] and quantum Monte Carlo [30, 31], the former method indicates a first-order transition, while the latter method suggests the existence of an intermediate gapless phase, as shown in Fig. 1a. Additionally, a fine-tuned transition between the TC and DS phases can be realized using deformed projected entangled pair states (PEPS) with some tuning parameters [32, 33, 34, 35]. Furthermore, a field-theoretic argument has been proposed for a continuous transition between the TC and DS phases in the presence of SU(2) symmetry [36]. Despite these efforts, how to concretely realize a direct continuous transition between the TC and DS phases using a generic Hamiltonian remains a long-standing open question. If such a continuous transition exists, the nature of the critical point, particularly how the incompatible anyons of the two $\mathbb{Z}_{2}$ topological orders reconcile at the critical point, is unclear.
In this letter, we confirm the existence of a direct continuous TC-DS transition and provide new insights into its nature. Rather than considering a model interpolating directly between fixed points of the TC and DS phases [29, 30, 31], we propose a perturbed $\mathbb{Z}_{4}$ quantum double (QD) model based on the observation that both phases can emerge from the $\mathbb{Z}_{4}$ topological order via anyon condensation [20, 22, 37]. To determine its phase diagram and the quantum criticality of the continuous TC-DS transition, we introduce a duality transformation that maps the perturbed $\mathbb{Z}_{4}$ QD model to a two-coupled (2+1)(2+1)D quantum Ising model and approximate its ground states with variationally optimized infinite PEPS (iPEPS) [38, 39] based on automatic differentiation [40].
Using the condensation order parameters and the area law coefficients of the Wilson loop expectation values, we identify that the continuous TC-DS transition belongs to the XY* universality class. Interestingly, as shown in Fig. 1b, the continuous TC-DS transition resembles the deconfined quantum critical point [41, 42]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, the perturbed $\mathbb{Z}_{4}$ QD model has two different 1-form symmetries, which are loop-like generalized symmetries. One 1-form symmetry is broken spontaneously in the TC phase, while the other one is broken spontaneously in the DS phase. Second, some anyons in the $\mathbb{Z}_{2}$ topological orders are fractionalized and become deconfined at the continuous TC-DS transition point. Our results suggest that at the critical point between topological phases that cannot be connected via anyon condensation, fractionalized anyons that are confined in both adjacent phases can emerge. This opens a path toward developing a general theoretical framework for topological phase transitions beyond anyon condensation.
A perturbed $\mathbb{Z}_{4}$ quantum double model. Because the model interpolating between the fixed-point Hamiltonians of the TC and DS phases [29, 30] is difficult to analyze analytically and simulate numerically, we alternatively propose a perturbed $\mathbb{Z}_{4}$ QD model to realize the topological phase transition between TC and DS phases  [^1] The $\mathbb{Z}_{4}$ QD model is also known called $\mathbb{Z}_{4}$ toric code model. In this work, “TC” refers exclusively to the $\mathbb{Z}_{2}$ toric code model, and we avoid calling $\mathcal{H}_{\text{QD}}$ the $\mathbb{Z}_{4}$ TC model to prevent confusion.. The fixed-point Hamiltonian for the $\mathbb{Z}_{4}$ QD phase on the square lattice is [4]:
$\mathcal{H}_{\text{QD}}=-\left(\sum_{v}A_{v}+\sum_{p}B_{p}+\text{h.c.}\right),$ (1)
[…]
According to anyon condensation theory [45, 46, 47, 48], the $\mathbb{Z}_{4}$ QD model can undergo phase transitions to the TC phase by condensing the charge $\pmb{e}^{2}$  [^2] Alternative, the phase transitions to the TC phase by condensing the flux $\pmb{m}^{2}$ by considering the Hamiltonian $\mathcal{H}_{\mathrm{QD}}-h_{x}\sum_{e}X_{e}^{2}$ with $h_{x}>0.657$. , and to the DS phase by condensing the dyon $\pmb{e}^{2}\pmb{m}^{2}$ [32, 33]. Microscopically, the transition to the TC phase can be realized by a Hamiltonian $\mathcal{H}_{\mathrm{QD}}-h_{z}\sum_{e}Z_{e}^{2}$ with $h_{z}\gtrsim 0.657$  [^3] The phase transition point can be determined by mapping the Hamiltonian $\mathcal{H}_{\mathrm{QD}}-h_{z}\sum_{e}Z_{e}^{2}$ to the transverse-field Ising model $H_{\text{Ising}}=-\sum_{\langle i,j\rangle}\sigma^{z}_{i}\sigma^{z}_{j}-h\sum_{i}\sigma^{x}_{i}$. This yields the critical point at $h_{z}=2/h_{c}^{\text{TFIM}}$, where $h_{c}^{\text{TFIM}}=3.044\,330(6)$ is the known critical value of the transverse-field Ising model [60].. In contrast, dyon condensation is less straightforward. Interestingly, the fixed-point DS model can be constructed using $\mathbb{Z}_{4}$ qudit operators [37]:
$\mathcal{H}_{\text{DS}}=-\sum_{p}\left(A_{v(p)}B_{p}+A_{v(p)}B_{p}^{\dagger}+\text{h.c.}\right)-\sum_{e}W^{2}_{e},$ (2)
where $v(p)$ denotes the vertex in the southwest of and adjacent to the plaquette $p$ (see Fig. 2a). The first term penalizes semionic dyon excitations $\pmb{e}^{n}\pmb{m}^{k}$ with odd $n,k$  [^4] Note that $A_{v}B_{p}A_{v}B_{p}^{\dagger}=A_{v}^{2}$ and $A_{v}B_{p}A_{v}^{\dagger}B_{p}=B_{p}^{2}$ are conserved quantities. Although the Hamiltonian used here differs from $\mathcal{H}^{\prime}_{\text{DS}}=-\sum_{\langle vp\rangle^{\prime}}(A_{v}B_{p}+B_{p}^{2}+\text{h.c.})-\sum_{e}W_{e}$ in Ref. [37], they share the same ground state.. The last term $W^{2}_{e}=X^{2}_{e}Z^{2}_{e^{\prime}}$ induces the condensation of the bosonic dyon $\pmb{e}^{2}\pmb{m}^{2}$, where $e^{\prime}$ denotes the nearest neighboring edge in the southwest of the edge $e$.
By combining $\mathcal{H}_{\text{QD}}$, $\mathcal{H}_{\text{DS}}$, and the perturbation terms $\sum_{e}X_{e}^{2}$ and $\sum_{e}Z_{e}^{2}$, we propose the following perturbed $\mathbb{Z}_{4}$ QD model to study topological phase transitions between the TC and DS phases:
$H_{\text{QD}}$ $=-\left(\sum_{v}A_{v}+\sum_{p}B_{p}+\text{h.c.}\right)-h_{x}\sum_{e}X_{e}^{2}-h_{z}\sum_{e}Z_{e}^{2}$
$\quad-\frac{1}{2}\sum_{p}\left(A_{v(p)}B_{p}+A_{v(p)}B_{p}^{\dagger}+\text{h.c.}\right)-h_{w}\sum_{e}W^{2}_{e}.$ (3)
The Hamiltonian possesses an extensive set of conserved quantities: $A^{2}_{v},H_{\text{QD}}]=0$ for all vertices, and $B^{2}_{p},H_{\text{QD}}]=0$ for all plaquettes. Multiplying these operators generates the so-called 1-form symmetries [52, 53, 54], represented by the non-local loop operators $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}W_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, where $C$ and $\hat{C}$ denote loops on the original and dual lattices, respectively. Since different topological orders can be understood as distinct patterns of higher-form symmetry breaking [55], the 1-form symmetries can be used to characterize different topological phases.
Figure 2: Perturbed $\mathbb{Z}_{4}$ QD model and duality transformation. (a) The multi-site local Hamiltonian terms of the perturbed $\mathbb{Z}_{4}$ QD model, $e$ of $W_{e}^{2}$ is the edge where the $X^{2}$ operator is located. (b) The perturbed $\mathbb{Z}_{4}$ QD model is mapped to a two-coupled Ising model, the blue (red) lattice is the primal (dual) lattice, and we shift the dual lattice to overlap with the primal lattice.
Condensation order parameters and area law coefficients. We can use non-local order parameters to detect spontaneous 1-form symmetry breaking, which implies different topological phases. A kind of such non-local order parameters includes $\left\langle\prod_{e\in S}Z_{e}^{2}\right\rangle$, $\left\langle\prod_{e\in\hat{S}}W_{e}^{2}\right\rangle$, and $\left\langle\prod_{e\in\hat{S}}X_{e}^{2}\right\rangle$ [56, 57], where $S$ ($\hat{S}$) denotes the open strings on the original (dual) lattices. They can be viewed as segments of closed loops $C$ ($\hat{C}$), as detailed in SM [44]. Because the anyons $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ are created at the endpoints of the string operators, and their non-zero expectation values imply the corresponding anyons condense, as summarized in Tab. 1, and we call them condensation order parameters.
Table 1: Condensation order parameters and area law coefficients. $S$ ($\hat{S}$) denotes a semi-infinite long string on the primal (dual) lattice, $C$ ($\hat{C}$) denote the a loop on the primal (dual) lattice, $R$ denotes the set of sites surround by the loop $C$ ($\hat{C}$) and $a_{X}$ is the area of the shape $X$. The column “Def. in $H_{\text{QD}}$” lists the definition of the condensation order parameters and area law coefficients in the $\mathbb{Z}_{4}$ QD model, while “Def. in $H_{\text{dual}}$” for the two-coupled Ising model. The columns $\mathbb{Z}_{4}$ QD”, “DS”, “TC” and “Trivial” show values of these quantities in the respective phases.
Label Def. in $H_{\text{QD}}$ Def. in $H_{\text{dual}}$ $\mathbb{Z}_{4}$ QD DS TC Trivial
Cond. $\langle\pmb{e}^{2}\rangle$ $\left|\left\langle\prod_{e\in S}Z^{2}_{e}\right\rangle\right|$ $\left|\left\langle\tau_{i}^{z}\right\rangle\right|$ 0 0 >0>0 >0>0
$\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}W^{2}_{e}\right\rangle\right|$ $\left|\left\langle\tau_{i}^{z}\sigma_{i}^{x}\right\rangle\right|$ 0 >0>0 0 >0>0
$\langle\pmb{m}^{2}\rangle$ $\left|\left\langle\prod_{e\in\hat{S}}X^{2}_{e}\right\rangle\right|$ $\left|\left\langle\sigma_{i}^{x}\right\rangle\right|$ 0 0 0 >0>0
Conf. $\alpha(\pmb{e})$ $\frac{\log\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle}{-a_{C}}$ $\frac{\log\langle\prod_{i\in R}\sigma_{i}^{z}\rangle}{-a_{R}}$ 0 >0>0 0 >0>0
$\alpha(\pmb{e}\pmb{m})$ $\frac{\log\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle}{-a_{\hat{C}}}$ $\frac{\log\langle\prod_{i\in R}\tau^{x}_{i}\sigma^{z}_{i}\rangle}{-a_{R}}$ 0 0 >0>0 >0>0
$\alpha(\pmb{m})$ $\frac{\log\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle}{-a_{\hat{C}}}$ $\frac{\log\langle\prod_{i\in R}\tau^{x}_{i}\rangle}{-a_{R}}$ 0 >0>0 >0>0 >0>0
Another kind of non-local order parameters is defined using Wilson loop operators $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}\right\rangle$, where $e^{\prime}$ denotes the edge immediately southwest of $e$, and the exponents $\gamma(e),\leavevmode\nobreak\ \hat{\gamma}(e)$ taking values $1$ or $3$ depend on the edge orientations [37, 44]. It is well known that the expectation values of the Wilson loop operators can detect the confinement of anyons [58, 59]. Let us focus on $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$, where the loop operator can be regarded as creating the pair $\pmb{e}$ and $\pmb{e}^{3}$ and annihilating them. When the expectation value satisfies $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle\propto\mathrm{e}^{-\alpha(\pmb{e})a_{C}}$ for a large loop $C$ that surrounds the area $a_{C}$, where $\alpha(\pmb{e})$ is the area law coefficient. When $\alpha(\pmb{e})>0$, it is the area law and it indicates $\pmb{e}$ is confined. If $\alpha(\pmb{e})=0$, it implies that $\left\langle\prod_{e\in C}Z^{\gamma(e)}_{e}\right\rangle$ decays at most as the perimeter of the loop $C$, i.e. the perimeter law, indicating that $\alpha(\pmb{e})$ is deconfined. Therefore, we can use the area law coefficient $\alpha(\pmb{e})$ to distinguish the deconfinement and confinement of $\pmb{e}$. A similar argument applies to $\left\langle\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right\rangle$ and $\left\langle\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right\rangle$, which is summarized in Tab. 1.
Figure 3: Phase diagram of the perturbed $\mathbb{Z}_{4}$ QD model on the $h_{x}=-1$ plane. The model has three tuning parameters: $h_{x}$, $h_{z}$ and $h_{w}$. The solid blue lines indicate the 3D Ising* transitions. The red star marks the multi-critical point, which belongs to the XY* universality class. Arrows (i) and (ii) denote the lines $h_{z}=h_{w}$ and $h_{w}=2.980-h_{z}$, respectively, both crossing the multi-critical point.
Mapping to a two-coupled (2+1)D quantum Ising model. The 1-form symmetries not only characterize different topological phases but also simplify the perturbed $\mathbb{Z}_{4}$ QD model in Eq. (Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases) via a duality transformation. It is well known that the $\mathbb{Z}_{2}$ lattice gauge theory with the exact 1-form Wilson or ’t Hooft loop symmetry can be mapped to the transverse-field Ising model [60] via the Wegner duality transformation [58, 11]. In the same spirit, we propose the following duality transformation for the perturbed $\mathbb{Z}_{4}$ QD model:
$A_{v}+A_{v}^{\dagger}\rightarrow 2\tau_{v}^{x},\,\,B_{p}+B_{p}^{\dagger}\rightarrow 2\sigma_{p}^{z},\,\,X^{2}_{e}\rightarrow\sigma^{x}_{p}\sigma^{x}_{p^{\prime}},\,\,Z^{2}_{e}\rightarrow\tau^{z}_{v}\tau^{z}_{v^{\prime}},$ (4)
where $\{\sigma_{p}^{\alpha}\}$ ($\{\tau_{v}^{\alpha}\}$) with $\alpha\in\{x,y,z\}$ is a set of Pauli matrices of $\mathbb{Z}_{2}$ qubits applying on the site of the dual (primal) lattice. As shown in Fig. 2b, by shifting the dual lattice such that it completely overlaps the primal lattice, and relabeling $v,v^{\prime}$ and $p,p^{\prime}$ using $i,j$, the perturbed $\mathbb{Z}_{4}$ QD model in Eq. (Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases) is mapped to a two-coupled (2+1)(2+1)D quantum Ising model:
$H_{\text{dual}}=$ $-2\sum_{i}\left(\tau^{x}_{i}+\sigma_{i}^{z}+\tau^{x}_{i}\sigma^{z}_{i}\right)-h_{x}\sum_{\langle ij\rangle}\sigma^{x}_{i}\sigma^{x}_{j}-h_{z}\sum_{\langle ij\rangle}\tau^{z}_{i}\tau_{j}^{z}$
$-h_{w}\sum_{\langle ij\rangle}\tau^{z}_{i}\tau_{j}^{z}\sigma^{x}_{i}\sigma^{x}_{j}.$ (5)
see detailed derivation in SM [44]. If the term $\sum_{i}\tau_{i}^{x}\sigma^{z}_{i}$ is absent and $h_{x}=h_{z}$, it becomes a standard (2+1)(2+1)D quantum Ashkin-Teller model. It is obvious that the phase diagram of $H_{\text{dual}}$ is symmetric about $h_{x}=h_{z}$, because the model is invariant under the exchange of $\sigma_{i}^{x}\leftrightarrow\tau_{i}^{z}$, $\sigma_{i}^{z}\leftrightarrow\tau_{i}^{x}$, and $h_{x}\leftrightarrow h_{z}$. Furthermore, its phase diagram is also symmetric about $h_{z}=h_{w}$, because $UH_{\text{dual}}(h_{x},h_{z},h_{w})U^{\dagger}=H_{\text{dual}}(h_{x},h_{w},h_{z})$, where the unitary transformation is $U=\prod_{i}\left(\frac{1+\sigma_{i}^{x}}{2}+\tau_{i}^{x}\frac{1-\sigma_{i}^{x}}{2}\right)$. Since duality transformations preserve the structure of phase diagrams, we can directly deduce the phase diagram of the $\mathbb{Z}_{4}$ QD model $H_{\text{QD}}$ from that of $H_{\text{dual}}$.
The non-local condensation order parameters and the Wilson loop expectation values of the perturbed $\mathbb{Z}_{4}$ QD model can be mapped to the local order parameters and the membrane-like disorder parameters of the two-coupled quantum Ising model [61, 62] (see SM [44] and Tab. 1). This allows us to extract condensation order parameters and the area law coefficients directly from the ground states of the two-coupled Ising model, which can be approximated using $C_{4v}$-symmetric iPEPS [39, 38, 40]. Compared to other methods such as quantum Monte Carlo, iPEPS is more suitable to evaluate non-local order parameters, and the method of extracting area law coefficients from iPEPS is proposed in Ref. [63]. By comparing the numerical results of the condensation order parameters and the area law coefficients with Tab. 1, we obtain the phase diagrams for various $h_{x}$ planes (see SM [44]), one of which is shown in Fig. 3. It includes the $\mathbb{Z}_{4}$ QD, TC, DS, and trivial phases, and their phase boundaries are described by the 3D Ising* universality class describing 1-form symmetry breaking transitions [64]. Next, we focus on the multi-critical point, especially the continuous TC–DS transition.
Figure 4: Condensation order parameters and area law coefficients near and at the continuous TC-DS transition point $(h_{x},h_{z},h_{w})\approx(-1.000,1.490,1.490)$. Different colors correspond to different order parameters, and the colors from light to dark correspond to the bond dimensions $(D,\chi)=(4,100),(5,100),(6,100),(7,100)$, where $D$ is the iPEPS bond dimension and $\chi$ is the bond dimension of the corner transfer matrix renormalization group method contracting the iPEPS [65, 66, 67]. The vertical dashed line marks the location of the multi-critical point ($h_{c}=1.490$). From the XY* universality class, we have three order parameter critical exponents $\beta_{\phi}$, $\beta_{\phi}^{\prime}$, $\beta_{-}$ and two correlation length critical exponents $\nu_{+}$ and $\nu_{-}$. (a) Area law coefficients along the line (i) in Fig. 3. (b) Condensation order parameters along the line (i). (c) Double-log plot of the area law coefficients in (a), where the slope of the straight lines is $2\nu_{+}$. (d) Double-log plot of the condensation order parameters in (b), where the slopes of the straight lines are $\beta_{\phi}$ and $\beta_{-}$, respectively. (e) Area law coefficients along the line (ii) in Fig. 3. (f) Condensation order parameters along the line (ii). (g) Double-log plot of the area law coefficients in (e), where the slope of the straight lines is $2\nu_{-}$. (h) Double-log plot of the condensation order parameters in (f), where the slope of the straight lines is $\beta_{\phi}^{\prime}$.
Continuous topological phase transition between the $\mathbb{Z}_{4}$ QD phase and the trivial phase. By tuning $h_{x}$ to negative values, we find that a continuous TC-DS phase transition becomes possible. We focus on the $h_{x}=-1$ plane (Fig. 3), where a continuous TC-DS transition emerges at the junction of four phases: the $\mathbb{Z}_{4}$ QD, TC, DS, and trivial phases. Since the phase diagram is symmetric about $h_{z}=h_{w}$ [line (i) in Fig. 3] and the continuous TC-DS transition must occur at a point on this line, we first analyze the condensation order parameters and area law coefficients along this line.
Fig. 4a shows the area law coefficients $\alpha(\pmb{e})$, $\alpha(\pmb{e}\pmb{m})$ and $\alpha(\pmb{m})$ are zero when $h_{w}=h_{z}<h_{c}\approx 1.490$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{n}$, $\pmb{e}^{n}\pmb{m}^{k}$ and $\pmb{m}^{k}$ ($n,k\in\{1,3\}$) are confined. Moreover, Fig. 4b shows the condensation order parameters $\langle\pmb{e}^{2}\rangle$, $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{m}^{2}\rangle$ are zero when $h_{w}=h_{z}<h_{c}$, but they are non-zero when $h_{w}=h_{z}>h_{c}$, implying $\pmb{e}^{2}$, $\pmb{e}^{2}\pmb{m}^{2}$ and $\pmb{m}^{2}$ condense. By comparing the numerical results in Figs. 4a and b with Tab. 1, we conclude the perturbed $\mathbb{Z}_{4}$ QD model is in the $\mathbb{Z}_{4}$ QD when $h_{w}=h_{z}<h_{c}$ and trivial phases when $h_{w}=h_{z}>h_{c}$.
Since the Ising* lines meet at the multi-critical point, a natural expectation of the effective theory describing the physics near and at the multi-critical point is the O(2) $\phi^{4}$ field theory, given by the Lagrangian $\mathcal{L}=\left(\partial\pmb{\phi}\right)^{2}/2+m^{2}\pmb{\phi}^{2}/2+g\pmb{\phi}^{4}/(4!)+\cdots$, where $\pmb{\phi}=(\phi_{0},\phi_{1})$ is a two-component order parameter field corresponding to the condensation order parameters $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ and $\langle\pmb{e}^{2}\rangle$. The multi-critical point corresponds to the Wilson-Fisher fixed point of the O(2) $\phi^{4}$ field theory, which is the 3D XY* conformal field theory [68, 69]. We can predict critical exponents of the condensation order parameters and the are law coefficients from the field theory [44]. In Fig. 4c, the area law coefficients match the correlation length critical exponent $\nu_{+}=1/(3-\Delta_{+})$, where $\Delta_{+}$ is the scaling dimension of the relevant field $\pmb{\phi}^{2}=\phi_{0}^{2}+\phi_{1}^{2}$ deriving to the gapped phases along $h_{w}=h_{z}$. Fig. 4d shows that the three condensation order parameters exhibit two distinct critical exponents $\beta_{\phi}=\Delta_{\phi}\nu_{+}$ and $\beta_{-}=\Delta_{-}\nu_{+}$, where $\Delta_{\phi}$ and $\Delta_{-}$ are the scaling dimensions of the fields $\phi_{0}$ (or $\phi_{1}$) and $\phi_{0}^{2}-\phi_{1}^{2}$, which correspond to $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ (or $\langle\pmb{e}^{2}\rangle$) and $\langle\pmb{m}^{2}\rangle$, separately. The numerical results imply that the multi-critical point belongs to the XY* universality class.
Continuous topological phase transition between the toric code and double semion phases. Finally, we investigate the path $h_{x}=-1,h_{w}=2h_{c}-h_{z}$ [line (ii) in Fig. 3], which traverses the TC phase, the DS phase, and the multi-critical point, and is orthogonal to $h_{w}=h_{z}$ [line (i)]. The area law coefficients in Fig. 4e indicate that $\pmb{e}$ and $\pmb{e}^{3}$ are confined when $h_{w}>h_{z}$ and $\pmb{e}^{n}\pmb{m}^{k}$ with $n,k\in\{1,3\}$ are confined when $h_{w}<h_{z}$, while $\pmb{m}$ and $\pmb{m}^{3}$ are confined as long as $h_{w}\neq h_{z}$. According to the condensation order parameter in Fig. 4f, $\pmb{e}^{2}\pmb{m}^{2}$ condenses when $h_{w}>h_{z}$, while $\pmb{e}^{2}$ condenses when $h_{w}<h_{z}$. By comparing the numerical results in Figs. 4e and f with Tab. 1, we know along the line (ii), it is the TC phase when $h_{z}>h_{w}$ and the DS phase when $h_{w}>h_{z}$.
As shown in Fig. 4g, we find that the area law coefficients exhibit a critical exponent consistent with the correlation length exponent $\nu_{-}=1/(3-\Delta_{-})$, where $\Delta_{-}$ is the scaling dimension of the relevant field $\phi_{0}^{2}-\phi_{1}^{2}$ deriving the multi-critical point to the TC or DS phases. Fig. 4h shows that the condensation order parameter exhibits the critical exponents $\beta_{\phi}^{\prime}=\Delta_{\phi}\nu_{-}$. The distinct critical exponents from the lines (i) and (ii) in Fig. 3 further confirm that the multi-critical point belongs to the XY* universality class, and the detailed field theory analysis is provided in SM [44].
Interestingly, the continuous TC-DS transition exhibits features that resemble deconfined quantum critical points [41, 42]—a direct continuous transition between two phases that spontaneously break different symmetries and exhibit fractionalized excitations—in the following aspects: First, from Fig. 4f, $\langle\pmb{e}^{2}\rangle=0$ in the DS phase while $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle=0$ in the TC phase, implying the gapped TC and DS phases spontaneously break two different 1-form symmetries, $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$, respectively. Second, the anyons $\pmb{m}$ and $\pmb{m}^{3}$ are confined in both the TC and DS phases, but they are deconfined at the continuous TC-DS transition point, indicated by the area law coefficient $\alpha(\pmb{m})$ shown in Fig. 4e; since $\pmb{m}\times\pmb{m}=\pmb{m}^{3}\times\pmb{m}^{3}=\pmb{m}^{2}$ is a well-defined anyon in both TC and DS phases, $\pmb{m}$ and $\pmb{m}^{3}$ are fractionalized anyons with respect to the $\mathbb{Z}_{2}$ topological order. A difference between the continuous TC-DS transition point and deconfined quantum critical points is that two broken 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$ and $\prod_{e\in\hat{C}}W_{e}^{2}$ do not have mixed anomalies.
Discussions and outlook. In this letter, we propose a perturbed $\mathbb{Z}_{4}$ QD model to study the topological phase transitions between the TC and DS phases, from which we find that the direct continuous phase transition exists between these two $\mathbb{Z}_{2}$ topologically ordered phases. We identify that the continuous TC-DS transition belongs to the XY* universality class, where deconfined $\mathbb{Z}_{4}$ anyons emerge. Our findings imply that if we consider a model without any explicit $\mathbb{Z}_{4}$ structure but exhibiting a continuous TC-DS transition, fractionalized anyons can emerge at the transition. We emphasize that the TC-DS transition is not a fine-tuned result since it is a continuous transition line in the plane defined by $h_{x}<0.8$ and $h_{z}=h_{w}$, and we further verified this conclusion through the analysis of O(2) field theory [44]. Since the TC and DS phases are dual to two different $\mathbb{Z}_{2}$ symmetry protected topological (SPT) phases in (2+1)(2+1)[70], the multi-critical point can also be interpreted as a topological critical point between two $\mathbb{Z}_{2}$ SPT phases. Our results shed light on constructing a general theoretical framework for continuous topological phase transitions beyond anyon condensation, i.e. considering emergent anyons to reconcile two incompatible topological orders.
Since the TC–DS transition can be simulated on a digital quantum computer [35], it would be interesting to experimentally realize this transition and measure the corresponding anyonic order parameters. In addition, one may consider adding more perturbations like $h_{x}^{\prime}\sum_{e}X_{e}+\text{h.c.}$, $h_{z}^{\prime}\sum_{e}Z_{e}+\text{h.c.}$, to the perturbed $\mathbb{Z}_{4}$ QD model in Eq. (Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases). These terms explicitly break the exact 1-form symmetries $\prod_{e\in C}Z_{e}^{2}$, $\prod_{e\in\hat{C}}X_{e}^{2}$, hence the duality transformation and non-local order parameters do not work anymore. In such cases, the Fredenhagen–Marcu order parameter can be used to detect anyon condensation [71, 72, 57], and an adiabatic approach can be used to investigate confinement [73]. Alternatively, the broken exact 1-form symmetry could be restored via quantum error correction [74].
Acknowledgements. The authors gratefully acknowledge valuable discussions with Lei Wang on numerical computations and support in providing computational resources. W.-T. Xu thanks R.-Z. Huang for helpful comments and acknowledges support from the Munich Quantum Valley, which is supported by the Bavarian state government with funds from the Hightech Agenda Bayern Plus.
Data availability. The code and data are available on Zenodo [75].
Supplemental Material: Continuous topological phase transition between $\mathbb{Z}_{2}$ topologically ordered phases
SI Anyon condensation theory of the $\mathbb{Z}_{4}$ quantum deouble model
In this section, we review the properties of anyons in the $\mathbb{Z}_{4}$ topological order, and apply the anyon condensation theory to the $\mathbb{Z}_{4}$ topological order. There are three types of anyons in the $\mathbb{Z}_{4}$ topological phase. The electric charges $\{\pmb{e},\pmb{e}^{2},\pmb{e}^{3}\}$ correspond to eigenvalues $\{\mathrm{i},-1,-\mathrm{i}\}$ of $A_{v}$, while the magnetic fluxes $\{\pmb{m},\pmb{m}^{2},\pmb{m}^{3}\}$ correspond to the same eigenvalues of $B_{p}$. Their composites, $\pmb{e}^{n}\pmb{m}^{k}$ ($n,k\in\{1,2,3\}$), are known as dyons. As an Abelian topological order, all anyons of the $\mathbb{Z}_{4}$ topological order are Abelian anyons with the quantum dimension $1$. The topological spins of the charges and the fluxes are $1$. The topological spins of the dyon $\pmb{e}^{n}\pmb{m}^{k}$ are determined by the mutual braiding phase factor $\mathrm{i}^{nk}$ between $\pmb{e}^{n}$ and $\pmb{m}^{k}$. The mutual braiding of the anyons $\pmb{e}^{n}\pmb{m}^{k}$ and $\pmb{e}^{q}\pmb{m}^{r}$ ($n,k,q,r\in\{0,1,2,3\}$) gives rise to the phase factor $\mathrm{i}^{nr-kq}$. The properties of these anyons are summarized in Tab. S1.
Table S1: Anyons of the $\mathbb{Z}_{4}$ topological order and condensation patterns in the TC and DS phases. The abbreviation “conf.” denotes confinement, and “cond.” denotes condensation.
Type Trivial Charge Flux Dyon
Anyon $\pmb{1}$ $\pmb{e}$ $\pmb{e}^{2}$ $\pmb{e}^{3}$ $\pmb{m}$ $\pmb{m}^{2}$ $\pmb{m}^{3}$ $\pmb{e}\pmb{m}$ $\pmb{e}\pmb{m}^{2}$ $\pmb{e}\pmb{m}^{3}$ $\pmb{e}^{2}\pmb{m}$ $\pmb{e}^{2}\pmb{m}^{2}$ $\pmb{e}^{2}\pmb{m}^{3}$ $\pmb{e}^{3}\pmb{m}$ $\pmb{e}^{3}\pmb{m}^{2}$ $\pmb{e}^{3}\pmb{m}^{3}$
Topological spin 1 1 1 1 1 1 1 i -1 -i -1 1 -1 -i -1 i
TC $\pmb{1}$ $\pmb{e}$ cond. $\pmb{e}$ conf. $\pmb{m}^{2}$ conf. conf. $\pmb{e}\pmb{m}^{2}$ conf. conf. $\pmb{m}^{2}$ conf. conf. $\pmb{e}\pmb{m}^{2}$ conf
DS $\pmb{1}$ conf. $\pmb{m}^{2}$ conf. conf. $\pmb{m}^{2}$ conf. $\pmb{e}\pmb{m}$ conf. $\pmb{e}\pmb{m}^{3}$ conf. cond. conf. $\pmb{e}\pmb{m}^{3}$ conf. $\pmb{e}\pmb{m}$
Then we consider anyon condensation in the $\mathbb{Z}_{4}$ topological order. The basic rules for Abelian anyon condensation are  [1, 2, 3, 4]:
  • Only bosonic anyons whose topological spins are $1$ can condense.
  • Anyons whose mutual braiding phase factors with the condensed anyons are not $1$ are confined.
  • Two anyons that differ by a condensed anyon via the fusion rules become identical.
Based on these rules, we can find that by condensing $\pmb{e}^{2}$, the remaining anyons are $\{\pmb{1},\pmb{e},\pmb{m}^{2},\pmb{e}\pmb{m}^{2}\}$, which are exactly the anyons of the toric code, as shown in Tab. S1. A different toric code can be obtained by condensing $\pmb{m}^{2}$ such that the remaining anyons are $\{\pmb{1},\pmb{m},\pmb{e}^{2},\pmb{e}^{2}\pmb{m}\}$. Moreover, there is a bosonic dyon $\pmb{e}^{2}\pmb{m}^{2}$ in the $\mathbb{Z}_{4}$ topological order. By condensing $\pmb{e}^{2}\pmb{m}^{2}$, the remaining anyons are $\{\pmb{1},\pmb{e}\pmb{m},\pmb{e}\pmb{m}^{3},\pmb{e}^{2}\}$, also in Tab. S1. These anyons are exactly the anyons $\{\pmb{1},\pmb{s},\bar{\pmb{s}},\pmb{b}=\pmb{s}\bar{\pmb{s}}\}$ of the DS, where $\pmb{s}$ denotes the semion.
SII derivation of the duality transformation
In this section, we present a detailed derivation of the duality transformation and discuss the underlying physics. The first step is to map the $\mathbb{Z}_{4}$ qudit operators on the edges of the lattice to the new $\mathbb{Z}_{4}$ qudit operators on the vertices and plaquettes (see Fig. 2b):
$A_{v}\rightarrow X_{v},\,\,B_{p}\rightarrow Z_{p},\,\,X_{e}\rightarrow X_{p(e)}X_{p^{\prime}(e)},\,\,Z_{e}\rightarrow Z_{v(e)}Z_{v^{\prime}(e)},$ (S1)
where $p(e)$ and $p^{\prime}(e)$ [$v(e)$ and $v^{\prime}(e)$] denote a pair of plaquettes [vertices] that are nearest neighboring to the edge $e$. This step takes advantage of conserved quantities of $H_{\text{QD}}$, and it can be checked that the set of operators $\{A_{v},B_{p},X_{e},Z_{e}|A^{2}_{v}=B_{p}^{2}=1\}$ and $\{X_{v},Z_{p},X_{p}X_{p^{\prime}},Z_{v}Z_{v^{\prime}}|X_{v}^{2}=Z_{p}^{2}=1\}$ satisfy the same algebra. Applying the transformation in Eq. (S1) to the perturbed $\mathbb{Z}_{4}$ QD model, we get:
$H_{\text{QD}}$ $=-\left[\sum_{v}X_{v}+\sum_{p}Z_{p}+\frac{1}{2}\sum_{p}\left(X_{v(p)}Z_{p}+X_{v(p)}Z^{\dagger}_{p}\right)+\text{h.c.}\right]-h_{x}\sum_{\langle pp^{\prime}\rangle}X^{2}_{p}X_{p^{\prime}}^{2}-h_{z}\sum_{\langle vv^{\prime}\rangle}Z^{2}_{v}Z_{v^{\prime}}^{2}-h_{w}\sum_{\langle pp^{\prime}\rangle}X^{2}_{p}X^{2}_{p^{\prime}}Z^{2}_{v(p)}Z^{2}_{v(p^{\prime})},$ (S2)
where $v(p)$ denotes the vertex $v$ adjacent to the plaquette $p$ and in the southwest direction of $p$, as shown in Fig. S1c. The above model can be further simplified by expressing the $\mathbb{Z}_{4}$ operators in terms of two sets of $\mathbb{Z}_{2}$ Pauli matrices $\{\sigma^{x},\sigma^{z}\}$ and $\{\tau^{x},\tau^{z}\}$:
$X_{v}+X_{v}^{\dagger}=(\mathbbm{1}_{v}+\sigma_{v}^{x})\otimes\tau_{v}^{x},\quad Z_{p}+Z_{p}^{\dagger}=\sigma_{p}^{z}\otimes(\mathbbm{1}_{p}+\tau_{p}^{z}),\quad Z_{v}^{2}=\mathbbm{1}_{v}\otimes\tau_{v}^{z},\quad X_{p}^{2}=\sigma_{p}^{x}\otimes\mathbbm{1}_{p},\quad X_{v}^{2}=\sigma_{v}^{x}\otimes\mathbbm{1}_{v},\quad Z_{p}^{2}=\mathbbm{1}_{p}\otimes\tau_{p}^{z}.$ (S3)
Due to the constraint $X_{v}^{2}=Z_{p}^{2}=1$, we have $\sigma_{v}^{x}=\tau_{p}^{z}=1$, which further simplifies the relations to:
$X_{v}+X_{v}^{\dagger}$ $=2\tau_{v}^{x},\quad Z_{p}+Z_{p}^{\dagger}=2\sigma_{p}^{z},\quad Z_{v}^{2}=\tau_{v}^{z},\quad X_{p}^{2}=\sigma_{p}^{x}.$ (S4)
Then, the Hamiltonian in Eq. (S2) can be rewritten in terms of these Pauli matrices:
$H_{\text{dual}}$ $=-2\left(\sum_{v}\tau^{x}_{v}+\sum_{p}\sigma_{p}^{z}+\sum_{p}\tau^{x}_{v(p)}\sigma^{z}_{p}\right)-h_{x}\sum_{\langle pp^{\prime}\rangle}\sigma^{x}_{p}\sigma^{x}_{p^{\prime}}-h_{z}\sum_{\langle vv^{\prime}\rangle}\tau^{z}_{v}\tau_{v^{\prime}}^{z}-h_{w}\sum_{\langle pp^{\prime}\rangle}\sigma^{x}_{p}\sigma^{x}_{p^{\prime}}\tau^{z}_{v(p)}\tau_{v(p^{\prime})}^{z}.$ (S5)
By shifting the dual lattice such that it overlaps with the primal lattice, and relabeling $v,v^{\prime}$ and $p,p^{\prime}$ using $i,j$, we obtain the two-coupled (2+1)(2+1)D quantum Ising model:
$H_{\text{dual}}=-2\sum_{i}\left(\tau^{x}_{i}+\sigma_{i}^{z}+\tau^{x}_{i}\sigma^{z}_{i}\right)-h_{x}\sum_{\langle ij\rangle}\sigma^{x}_{i}\sigma^{x}_{j}-h_{z}\sum_{\langle ij\rangle}\tau^{z}_{i}\tau_{j}^{z}-h_{w}\sum_{\langle ij\rangle}\tau^{z}_{i}\tau_{j}^{z}\sigma^{x}_{i}\sigma^{x}_{j}.$ (S6)
We validate the duality transformation by comparing the energy spectra of the perturbed $\mathbb{Z}_{4}$ QD model $H_{\text{QD}}$ and the dual model $H_{\text{dual}}$ using exact diagonalization. The agreement between the spectra of the two models confirms the correctness of the duality mapping.
In certain special limits, the two-coupled quantum Ising model in Eq. (S6) can be further simplified. When $h_{x}=h_{w}=0$, the degrees of freedom of $\{\sigma_{i}^{z}\}$ are fixed to $1$. The model reduces to a single (2+1)(2+1)D transverse-field Ising model (TFIM) : $H=-4\sum_{i}\tau_{i}^{x}-h_{z}\sum_{\langle ij\rangle}\tau_{i}^{z}\tau_{j}^{z}$. This yields the critical point at $h_{z}=4/h_{c}^{\text{TFIM}}$, where $h_{c}^{\text{TFIM}}=3.044\,330(6)$ is the known critical field of the TFIM [5]. When $h_{x}=0,\leavevmode\nobreak\ h_{w}\rightarrow+\infty$, there are constraints $\tau_{i}^{z}\tau_{j}^{z}=\sigma_{i}^{x}\sigma_{j}^{x}=1,\forall\langle i,j\rangle$. $H_{\text{dual}}$ reduces to another TFIM: $H=-2\sum_{i}s_{i}^{x}-h_{z}\sum_{\langle ij\rangle}s^{z}_{i}s^{z}_{j}$, where $s_{i}^{z}=\tau_{i}^{z}$ and $s_{i}^{x}=\tau_{i}^{x}\sigma_{i}^{z}$. Similarly, when $h_{x}=h_{z}=0$ and $h_{x}=0,\leavevmode\nobreak\ h_{z}\rightarrow+\infty$, $H_{\text{dual}}$ also reduces to the TFIM.
Figure S1: String and loop operators defining the condensation order parameters and the Wilson loop order parameters. (a) Sting operator creating two $\pmb{e}^{2}$ at $v$ and $v^{\prime}$. (b) Sting operator creating $\pmb{m}^{2}$ at $p$ and $p^{\prime}$. (c) Sting operator creating $\pmb{e}^{2}\pmb{m}^{2}$ at $p$ and $p^{\prime}$. (d) Wilson loop operator associated to $\pmb{e}$. (e) Wilson loop operator associated to $\pmb{m}$. (f) Wilson loop operator associated to $\pmb{e}\pmb{m}$.
Table S2: Symmetry breaking patterns in different topological phases of the perturbed $\mathbb{Z}_{4}$ QD model, and the corresponding phases and symmetries in the dual two-coupled quantum Ising model. The superscripts denote the form of symmetries.
Perturbed $\mathbb{Z}_{4}$ QD model $\mathbb{Z}_{4}$ QD TC DS Trivial
Ground state symmetry $\mathbb{Z}_{1}^{(1)}$ $\mathbb{Z}_{2}^{(1)}=\{\mathbbm{1},\prod_{e\in C}Z_{e}^{2}\}$ $\mathbb{Z}_{2}^{\prime(1)}=\{\mathbbm{1},\prod_{e\in\hat{C}}X_{e}^{2}Z_{e^{\prime}}^{2}\}$ $\mathbb{Z}_{2}^{(1)}\times\mathbb{Z}_{2}^{\prime(1)}$
Two-coupled quantum Ising model Paramagnet $\tau^{z}$ Ferromagnet $\sigma^{x}\tau^{z}$ Ferromagnet Baxter
Ground state symmetry $\mathbb{Z}_{2}^{(0)}\times\mathbb{Z}_{2}^{\prime(0)}$ $\mathbb{Z}_{2}^{(0)}=\{\mathbbm{1},\prod_{i}\sigma_{i}^{z}\}$ $\mathbb{Z}_{2}^{\prime(0)}=\{\mathbbm{1},\prod_{i}\tau_{i}^{x}\sigma_{i}^{z}\}$ $\mathbb{Z}_{1}^{(0)}$
Moreover, the $\mathbb{Z}_{4}$ QD model has a $\mathbb{Z}^{(1)}_{2}\times\mathbb{Z}^{\prime(1)}_{2}$ 1-form symmetry (the superscripts denote the form of the symmetry), which has different spontaneous symmetry breaking patterns in different topologically ordered phases, as shown in Tab. S2. After the duality transformation, the two-coupled quantum Ising model has a $\mathbb{Z}^{(0)}_{2}\times\mathbb{Z}^{\prime(0)}_{2}$ 0-form symmetry, and the spontaneous symmetry breaking patterns in different phases are also summarized in Tab. S2.
Using the duality transformation, the condensation order parameter and Wilson loop order parameters associated with the 1-form symmetries of the $\mathbb{Z}_{4}$ QD model can be mapped to the order and disorder parameters of the two-coupled Ising model with the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 0-form symmetries, respectively. In particular, using the relations $X^{2}_{e}\rightarrow\sigma^{x}_{p(e)}\sigma^{x}_{p^{\prime}(e)},\,\,Z^{2}_{e}\rightarrow\tau^{z}_{v(e)}\tau^{z}_{v^{\prime}(e)}$, where $p(e)$ [$v(e)$] denotes the plaquette [vertex] adjacent to the edge $e$, one can easily check that the string operators transform as:
$\prod_{e\in\hat{S}(p,p^{\prime})}X_{e}^{2}\rightarrow\sigma^{x}_{p}\sigma^{x}_{p^{\prime}},\quad\prod_{e\in S(v,v^{\prime})}Z_{e}^{2}\rightarrow\tau^{z}_{v}\tau^{z}_{v^{\prime}},\quad\prod_{e\in\hat{S}(p,p^{\prime})}W_{e}^{2}\longrightarrow\sigma^{x}_{p}\sigma^{x}_{p^{\prime}}\tau^{z}_{v(p)}\tau^{z}_{v(p^{\prime})},$ (S7)
where $S(v,v^{\prime})$ $\hat{S}(p,p^{\prime})$] denotes the string on the primal [dual] lattice whose two end points are at two vertices [plaquettes] $v$ [$p$] and $v^{\prime}$ $p^{\prime}$], see Figs. S1a-c. Furthermore, using the relations $A_{v}+A_{v}^{\dagger}\rightarrow 2\tau_{i}^{x},\,\,B_{p}+B_{p}^{\dagger}\rightarrow 2\sigma_{i}^{z}$, we can transform the Wilson loop operators of the $\mathbb{Z}_{4}$ QD model to the disorder operators of the dual model [6, 7]:
$\frac{1}{2}\left[\left(\prod_{e\in C}Z^{\gamma(e)}_{e}\right)+\text{h.c.}\right]$ $=\frac{1}{2}\left[\left(\prod_{p\in a_{C}}B_{p}\right)+\text{h.c.}\right]$ $\xrightarrow{\makebox[36.98866pt]{\smash{\scriptsize$B_{p}^{2}=1$}}}$ $\frac{1}{2^{|a_{C}|}}\prod_{p\in a_{C}}\left(B_{p}+B^{\dagger}_{p}\right)$ $\longrightarrow$ $\prod_{i\in R}\sigma^{z}_{i},$
$\frac{1}{2}\left[\left(\prod_{e\in\hat{C}}X^{\gamma(e)}_{e}\right)+\text{h.c.}\right]$ $=\frac{1}{2}\left[\left(\prod_{v\in a_{\hat{C}}}A_{v}\right)+\text{h.c.}\right]$ $\xrightarrow{\makebox[36.98866pt]{\smash{\scriptsize$A_{v}^{2}=1$}}}$ $\frac{1}{2^{|a_{\hat{C}}|}}\prod_{v\in a_{\hat{C}}}\left(A_{v}+A^{\dagger}_{v}\right)$ $\longrightarrow$ $\prod_{i\in R}\tau^{x}_{i},$
$\frac{1}{2}\left[\left(\prod_{e\in\hat{C}}X^{\hat{\gamma}(e)}_{e}Z^{\gamma(e^{\prime})}_{e^{\prime}}\right)+\text{h.c.}\right]$ $=\frac{1}{2}\left[\left(\prod_{p\in a_{C}}A_{v(p)}B_{p}\right)+\text{h.c.}\right]$ $\xrightarrow{\makebox[36.98866pt]{\smash{\scriptsize$A_{v}^{2}=B_{p}^{2}=1$}}}$ $\frac{1}{4^{|a_{C}|}}\prod_{p\in a_{C}}\left(A_{v(p)}B_{p}+A^{\dagger}_{v(p)}B_{p}+\text{h.c.}\right)$ $\longrightarrow$ $\prod_{i\in R}\tau^{x}_{i}\sigma^{z}_{i},$ (S8)
where $\gamma(e),\leavevmode\nobreak\ \hat{\gamma}(e)$ taking values $1$ or $3$ depends on the direction of the edge $e$, see Figs. S1d-f. The symbol $a_{C}$ ($a_{\hat{C}}$) is the set of plaquettes (vertices) surrounded by $C$ ($\hat{C}$). $R$ denotes the set of sites surround by the loop $C$ ($\hat{C}$) after shifting the dual lattice to the primal lattice. The derivation takes advantage of the fact that the ground states of the perturbed $\mathbb{Z}_{4}$ QD model satisfy $A_{v}^{2}=B_{p}^{2}=1$.
An interesting question regarding the duality transformation is: How can we explicitly construct the operator that implements the duality transformation? We can think about this question from the fact that two categories $\text{Vec}(\mathbb{Z}_{4})$ and $\text{Vec}^{\omega_{\text{II}}}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})$ are Morita equivalent [8, 9], where $\omega_{\text{II}}$ is a type-II cocycle of the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ group. Specifically, we can apply a finite depth unitary circuit (FDUC) to transform the perturbed $\mathbb{Z}_{4}$ QD model to a perturbed type-II cocycle twisted $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ QD model [10]. Then, via a quantum circuit whose depth grows at least linearly on the system size (LDQC) [11], we can transform the perturbed and type-II cocycle twisted $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ QD model to the perturbed but untwisted $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ QD model, whose topological order is simply two copies of the $\mathbb{Z}_{2}$ topological order. Finally, applying the tensor product of two Wegner duality (duality⊗2) transformations on the perturbed but untwisted $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ quantum double model, we arrive at the two-coupled Ising model. The effect of applying these transformations on the fixed point $\mathbb{Z}_{4}$ QD model can be summarized as:
$\leavevmode\nobreak\ \mbox{Vec}(\mathbb{Z}_{4})\xrightarrow{\text{FDQC}}\mbox{Vec}^{\omega_{II}}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\xrightarrow{\text{LDQC}}\mbox{Vec}(\mathbb{Z}_{2}\times\mathbb{Z}_{2})\xrightarrow{\text{duality}^{\otimes 2}}\mbox{ Paramagnet},$ (S9)
and the transformation for the perturbed $\mathbb{Z}_{4}$ QD model is the same one. Interestingly, it is known that the DS phase has the intrinsic sign problem [12, 13], but the duality transformation maps it to a paramagnetic phase without the intrinsic sign problem. The underlying reason is that the intrinsic sign problem has been removed by the LDQC. It would be interesting to explicitly work out the transformations shown in Eq. (S9) in the future.
SIII Supplemental numerical Results
In this work, the ground states of the two-coupled (2+1)(2+1)D quantum Ising model are approximated by $C_{4v}$-symmetric iPEPS [14, 15] with various bond dimensions $D$, using the automatic differentiation technique [16]. The contraction of the squared norm of iPEPS is performed approximately using the $C_{4v}$ symmetric corner transfer matrix renormalization group (CTMRG) algorithm [17, 18, 19], with environment bond dimensions $\chi$. After evaluating the order parameters and the disorder parameters of the two-coupled quantum Ising model, we map them back to the condensation order parameters and the area law coefficients of the perturbed $\mathbb{Z}_{4}$ QD model.
Figure S2: Phase diagrams of the perturbed $\mathbb{Z}_{4}$ QD model. The solid blue lines indicate continuous transitions which belong to the 3D Ising universality class. The dashed green line and dots represent first-order phase transitions. The red star marks the multi-critical point that to the XY* universality class. On the (a) $h_{x}=0.5$, (b) $h_{x}=0$, (c) $h_{x}=-1.0$ plane.
We show the phase diagrams of the perturbed $\mathbb{Z}_{4}$ QD model on various $h_{x}$ planes in Fig. S2, which exhibit three distinct types of structures. The phase diagrams for $0.1<h_{x}<1.0$ resemble Fig. S2a, where the TC and DS phases are not directly connected and the $\mathbb{Z}_{4}$ QD and trivial phases are separated by a first-order transition line. As $h_{x}$ decreases into the range $0.2<h_{x}<0.1$, the first-order line gradually shrinks into a single first-order transition point, such that the four gapped phases meet together, as depicted in Fig. S2b. For $0.8<h_{x}<-0.2$, numerical results suggest the meeting point of the four phases is associated with a continuous phase transition, but we find that its critical exponents may vary continuously with decreasing $h_{x}$. Previous studies of the 3D classical AT model suggest that there are several possibilities: weakly first-order [20], a wide crossover [21], or continuous but non-universal [22, 23]. So the meeting point in the range $0.8<h_{x}<-0.2$ needs further investigation. Moreover, in the range $1.0\leq h_{x}<-0.8$, the multi-critical point falls within the XY* universality class, and the shape of the phase diagram is simialr to that of $h_{x}=-1$ plane, as shown in Fig. S2c. Finally, when $h_{x}<-1$, corresponding to a strong antiferromagnetic coupling in the two-coupled Ising model, the perturbed $\mathbb{Z}_{4}$ QD model may exhibit a rich phase diagram, as a mixed-order phase is known to emerge in the 3D classical AT model [24, 25], which could imply the existence of a phase in the perturbed $\mathbb{Z}_{4}$ QD model where the $\mathbb{Z}_{2}$ topological orders of TC and DS coexist.
Figure S3: Condensation order parameters and area law coefficients along the line $(h_{x},h_{z})=(0.5,0.1)$ as a function of $h_{w}$. Different colors correspond to different order parameters, and the colors from light to dark correspond to the bond dimensions $(D,\chi)=(5,100),(5,160),(6,100),(7,100)$. The vertical black dashed line corresponds to the location of the critical point $h_{w}\approx 1.292$, denoted by $h_{c}^{\text{Ising}}$. The 3D Ising critical exponents for order parameter and correlation length are approximately $\beta\approx 0.326$ and $\nu\approx 0.630$, respectively. (a) Area law coefficients. (b) Condensation order parameters. (c) Double-log plot of the area law coefficients $\alpha(\pmb{e})$ and $\alpha(\pmb{m})$, where the slope of the straight line is $2\nu$. (d) Double-log plot of the condensation order parameter $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$, where the slop of the straight line is $\beta$.
We consider the line $(h_{x},h_{z})=(0.5,0.1)$ with varying $h_{w}$, which crosses the $\mathbb{Z}_{4}$ QD and DS phases in the perturbed $\mathbb{Z}_{4}$ QD model. As shown in Fig. S3, we find that all the physical quantities converge well for iPEPS bond dimension $D=5$, and further increasing the CTMRG environment bond dimension $\chi$ has a negligible effect. Figure S3a shows that when $h_{w}>h_{c}^{\text{Ising}}$, the area law coefficients $\alpha(\pmb{e})>0$ and $\alpha(\pmb{m})>0$ but $\alpha(\pmb{e}\pmb{m})=0$, implying that the charges $\pmb{e}$ and $\pmb{e}^{3}$ and fluxes $\pmb{m}$ and $\pmb{m}^{3}$ are confined while the semionic dyon $\pmb{e}^{n}\pmb{m}^{k}$ with $n,k\in\{1,3\}$ remains deconfined. In Fig. S3b, the bosonic dyon $\pmb{e}^{2}\pmb{m}^{2}$ condenses ($\langle\pmb{e}^{2}\pmb{m}^{2}\rangle>0$) when $h_{w}>h_{c}^{\text{Ising}}$, while the charge $\pmb{e}^{2}$ and flux $\pmb{m}^{2}$ do not condense ($\langle\pmb{e}^{2}\rangle=\langle\pmb{m}^{2}\rangle=0$). Thus, the surviving anyons precisely match those of the DS topological order, see Sec. SI. In the two-coupled quantum Ising model, the above phase transition is a conventional 0-form $\mathbb{Z}_{2}$ symmetry breaking transition, where $\prod_{i}\sigma_{i}^{z}$ is spontaneously broken when $h_{w}>h_{c}^{\text{Ising}}$, so it belongs to the 3D Ising universality class. In the perturbed $\mathbb{Z}_{4}$ QD model, the corresponding transition involves the breaking of a $\mathbb{Z}_{2}$ 1-form symmetry, with $\prod_{e\in\hat{C}}W^{2}_{e}$ is spontaneously broken in the $\mathbb{Z}_{4}$ QD phase and preserved in the DS phase. This is known as the Ising* transition [26], which has the same critical exponents as the Ising transition. Hence, in the critical regime, the condensation order parameter satisfies $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle\propto(h_{w}-h_{c}^{\text{Ising}})^{\beta}$ and the area law coefficients satisfies $\alpha(\pmb{e})\propto\alpha(\pmb{m})\propto(h_{w}-h_{c}^{\text{Ising}})^{2\nu}$, where $\beta=0.326418(2)$ and $\nu=0.629970(4)$ [27]. As shown in Figs. S3c and d, the double-log plots of the numerical condensation order parameters and the area law coefficients indicate that the numerical results are consistent with the theoretical analysis.
Figure S4: Condensation order parameters and the area law coefficients along the line $h_{x}=-1.000$, $h_{w}=h_{z}$. A first-order transition occurs at $h_{w}=h_{z}\approx 0.980$. Different colors correspond to different anyons, and the colors from light to dark correspond to the bond dimensions $(D,\chi)=(4,100),(4,160),(5,100),(6,100)$. (a) Condensation order parameters. (b) Area law coefficients.
When $h_{x}=0.5$, there is no direct transition between the TC and DS phases, and the $\mathbb{Z}_{4}$ QD and trivial phases are separated by a first-order transition line, see Fig. S2a. We perform simulations along the line where $h_{w}=h_{z}$ as an example, as illustrated in Fig. S4. Near the first-order transition point $(h_{x},h_{z},h_{w})\approx(0.500,0.980,0.980)$, the condensation order parameters and the area law coefficients display discontinuities, which indicate a first-order phase transition. Furthermore, we find that all expectation values converge as the iPEPS bond dimension $D\geq 5$, and confirm that these quantities remain unchanged with increasing CTMRG environment bond dimension $\chi$.
SIV Analysis the critical exponent at the TC-DS transition point using $O(2)$ field theory
In this section, we show how to predict the critical exponents of the condensation order parameters and the area law coefficients at the TC-DS transition point, i.e. the multi-critical point where four phases meet. A natural candidate for effective Ginzburg-Landau field theory at the multi-critical point of the two-coupled Ising model is the O(2) $\phi^{4}$ model described by the Lagrangian:
$\mathcal{L}=\left(\partial\pmb{\phi}\right)^{2}/2+m^{2}\pmb{\phi}^{2}/2+g\pmb{\phi}^{4}/(4!)+\cdots,$ (S10)
where $m$ is the mass and $g$ is the coupling strength, and $\pmb{\phi}=(\phi_{0},\phi_{1})$ is a two-component order parameter field corresponding respectively to the two local order parameters $\langle\tau^{z}\sigma^{x}\rangle$ and $\langle\tau^{z}\rangle$. The Wilson-Fisher fixed point of the O(2) $\phi^{4}$ model belongs to the XY universality class [28, 29]. However, when $\mathcal{L}$ describes the physics near the multi-critical point of the perturbed $\mathbb{Z}_{4}$ QD model, $\pmb{\phi}=(\phi_{0},\phi_{1})$ corresponds to the two condensation order parameters $\langle\prod_{e\in\tilde{C}}W^{2}_{e}\rangle$ and $\langle\prod_{e\in C}Z^{2}_{e}\rangle$, and the Wilson-Fisher fixed point of $\mathcal{L}$ belongs to so-called XY* universality class. The starred and non-starred versions of the Ginzburg-Landau-Wilson (GLW) theory differ in the following aspects [26, 32]: First, in the starred version $\phi_{i}$ can only be created in pairs on a torus just like anyons; Second, $\phi_{i}$ and $\phi_{i}$ are indistinguishable in the started version, which is so-called orbifording, so that both periodic and anti-periodic boundary conditions have to be considered; Third, the starred version GLW theory describes the higher-form symmetry breaking transition, while the non-stared version describing the unusual symmetry breaking transition. Nevertheless, the scaling dimensions of certain fields in the XY and XY* universality classes are the same.
Table S3: Classifying fields in the O(2) $\phi^{4}$ model using irreps of the O(2), and the scaling dimensions of the fields are also shown.
field irrep. scaling dimension [30]
$\phi^{2}_{0}+\phi^{2}_{1}$ $\pmb{0}$ 1.51136(22)
$(\phi_{0},\phi_{1})$ $\pmb{1}$ 0.519088(22)
$(\phi^{2}_{0}-\phi^{2}_{1},\phi_{0}\phi_{1})$ $\pmb{2}$ 1.23629(11)
We can classify the fields in the O(2) $\phi^{4}$ model in terms of irreps of O(2) group, which includes the one-dimensional trivial irrep $\pmb{0}$, the one-dimensional sign irrep $\pmb{0}^{-}$ and two-dimensional irreps $\pmb{q}=q\oplus(-q)$, where the integer $q$ are the O(2) charges. Some fields transform under the irreps of O(2) and their scaling dimensions at the Wilson-Fisher fixed point are shown in Tab. S3.
Let us first consider the $\mathbb{Z}_{4}$ QD-trivial phase transition along the line (i): $h_{w}=h_{z},\leavevmode\nobreak\ h_{x}=-1$. In the corresponding two-coupled Ising model, the lattice operator that derives this transition is $\sum_{\langle i,j\rangle}\left(h_{w}\tau_{i}^{z}\tau_{j}^{z}\sigma_{i}^{x}\sigma_{j}^{x}+h_{z}\tau_{i}^{z}\tau_{j}^{z}\right)$, which is symmetric under the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry. So, we expect that the corresponding field is $\phi_{0}^{2}+\phi_{1}^{2}$, because it is also invariant under the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ symmetry: $\phi_{0}\leftrightarrow-\phi_{0}$ and $\phi_{1}\leftrightarrow-\phi_{1}$. Since the scaling dimension of $\phi_{0}^{2}+\phi_{1}^{2}$ is $\Delta_{\pmb{0}}=1.51136(22)$, the correlation length critical exponent is $\nu_{\pmb{0}}=1/(3-\Delta_{\pmb{0}})=0.67175(1)$. Because the quantity $\alpha[\cdot]\xi^{2}$ is dimensionless [31], the critical exponent of all area law coefficients is $2\nu_{\pmb{0}}$. Moreover, considering that the local order parameters $\langle\tau_{i}^{z}\rangle$ and $\langle\tau_{i}^{z}\sigma_{i}^{x}\rangle$ correspond to the fields $\phi_{0}$, $\phi_{1}$ with the scaling dimension $\Delta_{\pmb{1}}=0.519088(22)$ , we know $\langle\sigma_{i}^{x}\rangle$ corresponds to the field $\phi_{0}\phi_{1}$ with the scaling dimension $\Delta_{\pmb{2}}=1.23629(11)$. So the critical exponent of $\langle\tau_{i}^{z}\rangle$ and $\langle\tau_{i}^{z}\sigma_{i}^{x}\rangle$ is $\beta_{-}=\Delta_{\pmb{1}}\nu_{\pmb{0}}=0.34870(7)$, and the critical exponent of $\langle\sigma_{i}^{x}\rangle$ is $\beta_{\phi}=\Delta_{\pmb{2}}\nu_{\pmb{0}}=0.83048(2)$. Mapping the local order parameters of the two-coupled quantum Ising model back to the condensation order parameters of the perturbed $\mathbb{Z}_{4}$ QD model, we know the critical exponents of $\langle\pmb{e}^{2}\rangle$ and $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ are $\beta_{-}$, and the critical exponent of $\langle\pmb{m}^{2}\rangle$ is $\beta_{\phi}$.
Then, we consider the transition between the TC and DS phases along the line (ii): $h_{x}=-1,\leavevmode\nobreak\ h_{w}=2h_{c}-h_{z}$. When perturbing away from the critical point along this line, the $\mathbb{Z}_{2}$ symmetry exchanging $h_{z}$ and $h_{w}$ is broken. This corresponds to breaking the $\mathbb{Z}_{2}$ symmetry $\phi_{0}\leftrightarrow\phi_{1}$ in the field theory. Thus, the field $\phi_{0}^{2}-\phi_{1}^{2}$ with the scaling dimension is $\Delta_{\pmb{2}}=1.23629(11)$ is a candidate relevant field perturbing from the multi-critical point towards TC or DS phases. Therefore, the correlation length critical exponent $\nu_{\pmb{2}}=1/(3-\Delta_{\pmb{2}})=0.56699(4)$ determining the critical exponent of the area law coefficients is different from that of the transition between the $\mathbb{Z}_{4}$ QD and the trivial phases. Because in the two-coupled quantum Ising model, $\langle\tau_{i}^{z}\sigma_{i}^{x}\rangle$ and $\langle\tau_{i}^{z}\rangle$ correspond to $\phi_{0}$ and $\phi_{1}$ with the scaling dimension $\Delta_{\pmb{1}}=0.519088(22)$, the critical exponent of $\langle\tau_{i}^{z}\sigma_{i}^{x}\rangle$ is $\beta_{\phi}^{\prime}=\Delta_{\pmb{1}}\nu_{\pmb{2}}=0.519088(22)\times 0.56699(4)=0.29432(3)$. Mapping the local order parameters of the two-coupled Ising model back to the condensation order parameters of the perturbed $\mathbb{Z}_{4}$ QD model, we know the critical exponents of $\langle\pmb{e}^{2}\rangle$ and $\langle\pmb{e}^{2}\pmb{m}^{2}\rangle$ are $\beta^{\prime}_{\phi}$.
Next, we discuss the stability of the continuous TC-DS transition under varying $h_{x}$. The term $h_{x}\sum_{\langle i,j\rangle}\sigma^{x}_{i}\sigma^{x}_{j}$ is invariant under the $\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ spin-flip symmetry and the $\mathbb{Z}_{2}$ symmetry exchanging $h_{z}$ and $h_{w}$. Considering $\sigma^{x}_{i}$ corresponds to the field $\phi_{0}^{2}\phi_{1}^{2}$ [32], we expect that $\sigma^{x}_{i}\sigma^{x}_{j}$ corresponds to the field $\phi_{0}^{2}\phi_{1}^{2}$, which is a basis of the O(2) irrep $\pmb{4}$ with a scaling dimension $3.14(2)>3$ [30], so it is an irrelevant perturbation and the continuous TC-DS transition point is stable under varying $h_{x}$.
Finally, since it is also proposed that the O(2) $\phi^{4}$ theory can describe the topological multi-critical point of the standard $\mathbb{Z}_{2}$ TC model in a tilted field [33], let us compare the multi-critical point of the $\mathbb{Z}_{2}$ TC model in a tilted field and that of the perturbed $\mathbb{Z}_{4}$ QD model. The $\mathbb{Z}_{2}$ TC model in a tilted field has phase transitions from the TC phase to the trivial phase by condensing either $\mathbb{Z}_{2}$ charge excitations or $\mathbb{Z}_{2}$ flux excitations. Along the self-dual line, the duality symmetry swaps the $\mathbb{Z}_{2}$ charge and flux excitations, and there is a topological multi-critical point separating the TC phase and a first-order line where the duality symmetry is broken spontaneously [34, 35, 36]. Although it is proposed that the O(2) $\phi^{4}$ model can describe the topological multi-critical point of the $\mathbb{Z}_{2}$ TC model in a tilted field [33], the non-trivial mutual statistics between the $\mathbb{Z}_{2}$ charge and the flux excitations of the TC model are not included in the O(2) $\phi^{4}$ model because the two fields $\phi_{0}$ and $\phi_{1}$ commute with each other. Therefore, it is argued that the multi-critical point of the standard TC might not be the XY* universality class [37]. In contrast, there is no such problem at the multi-critical point of the perturbed $\mathbb{Z}_{4}$ QD model, because the mutual statistics between $\pmb{e}^{2}$ and $\pmb{e}^{2}\pmb{m}^{2}$ are bosonic.